Determine algebraically whether f(x)=x^3-x^2+1 is odd,even,neither.
Answer by sachi(392)

Determine algebraically whether f(x)=x^3-x^2+1 is odd,even,neither.

ans:-
if f(x)=f(-x) then the function is even
if -f(x)=f(-x) then the function is odd
so f(-x)=(-x)^3-(-x)^2+1=-x^3-x^2+1=-(x^3+x^2-1)=/-f(x)
so the function is neither even nor odd ans